OpenStudy (anonymous):

what is (2/8)±(squareroot (76 i )) /8 simplified???

OpenStudy (anonymous):

You sure you mean to have square root i there? It's not that hard to find, but a little uncommon.

OpenStudy (anonymous):

Im sorry, its kind of hard to type. Its (2/8) plus or minus ((square root 76)/8 )*i

OpenStudy (anonymous):

The question im trying to solve overall is: Find all solutions to the equation 4x^2 - 2x +5 = 0 and express them in the form a + bi.

OpenStudy (anonymous):

\[\frac{2}{8} \pm \frac{i \sqrt{76}}{8}\] ?

OpenStudy (anonymous):

What you just typed is basically what I was trying to write. Thats what I got when I put the equation above into the quadratic formula, but I think it can be simplified. Either that or I did something wrong

OpenStudy (anonymous):

It can be simplified (a little), but the answers are right. \[\frac{2}{8} = \frac{1}{4}\] \[\sqrt{76} = 2\sqrt{9} \implies \frac{\sqrt{76}}{8} = \frac{\sqrt{19}}{4}\] So the answers are \[\frac{1}{4} \pm \frac{\sqrt{19}}{4}i\]

OpenStudy (anonymous):

sqrt 76 = 2 sqrt (9) is a typo in the middle; should be 2 sqrt(19)

OpenStudy (anonymous):

What you just typed is basically what I was trying to write. Thats what I got when I put the equation above into the quadratic formula, but I think it can be simplified. Either that or I did something wrong

OpenStudy (anonymous):

Thank you I totally forgot I could break the square root apart!